hi my daughter has come home with and algebra question and I can't remember how to do it
question is a+b/c =7 and a-b/c=1 what is b?
Sorry i don't know where the underscore is but both formulas are over c.
thanks

Answer by richwmiller(17219)

not underscore but forward slash under the ?
We can't get a number value for b since we have three unknowns(variables) and only two equations.
We can only find b in terms of a or in terms of c
(a+b)/c=7
(a-c)/c=1
multiply both equations both sides by c
a+b=7c (1)
a-c=c
a=2c
c=a/2
substitute a/2 for c in equation (1)
a+b=7*a/2
multiply by 2
2a+2b=7a
2b=5a
b=5a/2